What fraction of an inch is one centimeter
1
Expert's answer
2019-12-19T10:09:50-0500

1 in=2.54 cm.

The corresponding fraction is

"\frac{1}{2.54}=\frac{100}{254}=\frac{50}{127}\approx\frac{25}{64}"